# Needle: A Generative-AI Powered Monte Carlo Method for Answering Complex Natural Language Queries on Multi-modal Data

### v1

@@ 5.1. Rankings Aggregation and Embedder Trust Mechanism
5.1.1. Aggregation
Given the ranked lists from all the embedders and the topic $t$ of the query $\varphi$, the objective is to aggregate these lists into a single final ranking $R_{\varphi}$ by incorporating the topic-specific reliability weights $w^{t}_{i}$.
Let $I_{h}\in\mathcal{D}$ be a tuple returned by at least one of the embedders. Let $rank(I_{h},R^{i}_{j})$ be the rank assigned to $I_{h}$ according to the list $R^{i}_{j}$. Let $\mathsf{S}:[k]\rightarrow[0,1]$ be a monotonically decreasing function that specifies the importance of each rank position in the top-$k$. In particular, while being agnostic to the choice of $S(.)$, following the literature (Cummins, 2012; Arampatzis and Robertson, 2011; Faria et al., 2010), we adopt the following position-importance function in Needle:
@@ 5.1.1. Aggregation
(4) $score(I_{h})=\sum_{i:\mathcal{E}^{i}\in\mathbf{E}}w^{t}_{i}\cdot\mathsf{S}(rank(I_{h},R^{i}_{j}))$
@@
5.1.2. Dynamic Reliability-weight Adjustment
Suppose after showing the top-$k$ results $R_{\varphi}$ to the user, we receive feedback from him/her, which is a subset of results marked as irrelevant to the query $\varphi$  [^3] We acknowledge that such feedback may not commonly be available after the system deployment. However, crowdsourced feedback on a small set of topic-based benchmarks can be employed to calculate these weights before deploying a new embedder to our system..
@@ 5.1.2. Dynamic Reliability-weight Adjustment
Using this feedback, we update the reliability weights of the embedders for each topic using the Multiplicative Weight Update Method (MWUM) (Arora et al., 2012). Let $t=t(\varphi)$ be the topic of the query $\varphi$. Then the partial loss of the embedder $\mathcal{E}^{i}$ for a topic $t$ based on $\varphi$ is computed as,
(5) $\Delta loss_{\varphi}(t,i)=\sum_{j:\bar{\mathbf{g}}_{j}\in\mathbf{G}}\sum_{I^{\prime}_{h}\in\mathcal{T}}\mathsf{S}(rank(I^{\prime}_{h},R^{i}_{j}))$
@@
5.2. Anomaly Detection in Query Images
@@
5.2.1. Dimensionality Reduction using UMAP (McInnes et al., 2018)
Given the high-dimensional nature of the embeddings, we first apply dimensionality reduction to project the embeddings into a lower-dimensional space. For each embedder $\mathcal{E}^{i}$, we reduce the dimensionality of the embeddings $\{\vec{v}^{i}_{1},\vec{v}^{i}_{2},\dots,\vec{v}^{i}_{m}\}$ to a lower dimension $d$ using UMAP (Uniform Manifold Approximation and Projection) (McInnes et al., 2018).
@@
5.2.2. Outlier Detection using LOF (Breunig et al., 2000)
After dimensionality reduction, we apply the Local Outlier Factor (LOF)(Breunig et al., 2000) algorithm to identify outliers. LOF assigns an outlier score to each data point based on the density of its neighborhood. The LOF score for image $\bar{\mathbf{g}}_{j}$ with respect to embedder $\mathcal{E}^{i}$ is defined as:
@@
5.2.3. Aggregation of Outlier Scores
The final outlier score of an guide image $\bar{\mathbf{g}}_{j}$ is computed by aggregating the individual scores, weighted with the reliability weight of each embedder:
@@ 5.2.3. Aggregation of Outlier Scores
5.2.4. Anomaly Detection Threshold
In the end, if the outlier score $S(\bar{\mathbf{g}}_{j})$ of an image $\bar{\mathbf{g}}_{j}$ exceeds a given threshold $\tau$, the image is flagged as an anomaly:

### v2

@@ C.1 Rankings Aggregation and Embedder Trust Mechanism
Aggregation: Given the ranked lists from all the embedders and the topic $t$ of the query $\varphi$, the objective is to aggregate these lists into a single final ranking $R_{\varphi}$ by incorporating the topic-specific reliability weights $w^{t}_{i}$.
Let $I_{h}\in\mathcal{D}$ be a tuple returned by at least one of the embedders. Let $rank(I_{h},R^{i}_{j})$ be the rank assigned to $I_{h}$ according to the list $R^{i}_{j}$. Let $\mathsf{S}:[k]\rightarrow[0,1]$ be a monotonically decreasing function that specifies the importance of each rank position in the top-$k$. In particular, while being agnostic to the choice of $S(.)$, following the literature [13, 4, 19], we adopt the following position-importance function in Needle:
@@
$score(I_{h})=\sum_{i:\mathcal{E}^{i}\in\mathbf{E}}w^{t}_{i}\cdot\mathsf{S}(rank(I_{h},R^{i}_{j}))$ (5)
@@
Dynamic Reliability-weight Adjustment: Suppose after showing the top-$k$ results $R_{\varphi}$ to the user, we receive feedback from him/her, which is a subset of results marked as irrelevant to the query $\varphi$  [^14] We acknowledge that such feedback may not commonly be available after the system deployment. However, crowdsourced feedback on a small set of topic-based benchmarks can be employed to calculate these weights before deploying a new embedder to our system..
@@
Using this feedback, we update the reliability weights of the embedders for each topic using the Multiplicative Weight Update Method (MWUM) [5]. Let $t=t(\varphi)$ be the topic of the query $\varphi$. Then the partial loss of the embedder $\mathcal{E}^{i}$ for a topic $t$ based on $\varphi$ is computed as,
$\Delta loss_{\varphi}(t,i)=\sum_{j:\bar{\mathbf{g}}_{j}\in\mathbf{G}}\sum_{I^{\prime}_{h}\in\mathcal{T}}\mathsf{S}(rank(I^{\prime}_{h},R^{i}_{j}))$ (6)
@@
C.2 Anomaly Detection in Query Images
@@
Dimensionality Reduction using UMAP [45]: Given the high-dimensional nature of the embeddings, we first apply dimensionality reduction to project the embeddings into a lower-dimensional space. For each embedder $\mathcal{E}^{i}$, we reduce the dimensionality of the embeddings $\{\vec{v}^{i}_{1},\vec{v}^{i}_{2},\dots,\vec{v}^{i}_{m}\}$ to a lower dimension $d$ using UMAP (Uniform Manifold Approximation and Projection) [45].
@@
Outlier Detection using LOF [8] : After dimensionality reduction, we apply the Local Outlier Factor (LOF)[8] algorithm to identify outliers. LOF assigns an outlier score to each data point based on the density of its neighborhood. The LOF score for image $\bar{\mathbf{g}}_{j}$ with respect to embedder $\mathcal{E}^{i}$ is defined as:
@@
Aggregation of Outlier Scores: The final outlier score of an guide image $\bar{\mathbf{g}}_{j}$ is computed by aggregating the individual scores, weighted with the reliability weight of each embedder:
@@
Anomaly Detection Threshold: In the end, if the outlier score $S(\bar{\mathbf{g}}_{j})$ of an image $\bar{\mathbf{g}}_{j}$ exceeds a given threshold $\tau$, the image is flagged as an anomaly: